# The Geometry of Intelligence A Functional–Topological Framework for Perception and Representation

### v2

@@ 1 Introduction
Throughout this work, $\mathcal{M}$ denotes the set of realized signals generated by the physical system—the observed perceptual manifold—rather than the full sensory space. When the governing equations are unknown, both the manifold structure and its radius must be inferred directly from the stream of observations, leading naturally to a self-supervised formulation. This view is consistent with contemporary approaches to autonomous intelligence lecun2022path, which emphasize that learning arises from discovering the set of admissible representations produced by the world rather than from external supervision.
@@
    We introduce a topological framework in which real-world perceptual sets are modeled as compact subsets of $C^{0}$ with finite Hausdorff radius and stable invariants rudin1991functional; royden2010real.
@@
    We show that deterministic physical processes induce continuous perceptual functionals that can be approximated by universal function approximators cybenko1989approximation; hornik1991approximation.
@@ 2.2 Compactness of the perceptual set
This result follows directly from the Arzelà–Ascoli theorem royden2010real; rudin1991functional. Compactness implies that the system cannot produce arbitrary variability: all realized signals lie inside a bounded, closed, finite-variability geometric region in function space.
@@ 3.2 Universal approximation of perceptual mappings
Proof: See Appendix A. Follows from the Universal Approximation Theorem cybenko1989approximation; hornik1991approximation and compactness of $\mathcal{M}$.
@@ 4.1 The perceptual radius
Using the Hausdorff metric $d_{H}$ edgar2008measure, the perceptual radius is defined as:
@@ 5.3 Distance metric and Hausdorff evaluation
In finite dimensions, all norms are equivalent (rudin1991functional, Ch. 1), so compactness, convergence, and the existence of a finite perceptual radius are preserved under cosine distance, although the numerical value of the radius may differ between metrics.
@@ 5.6.1 Railway Point Machine Current Traces
We use the public Chinese Railway Point Machine dataset li2020pointmachine; CRPM_dataset_url. The signals exhibit a characteristic deterministic structure (inrush, plateau, closure peak), ideal for testing compactness and Hausdorff radius.
@@ 5.6.2 NASA Battery Aging Dataset
We use the NASA Ames Battery Dataset NASA_Battery_Repository; saha2007prognostics; saha2011modeling. Battery discharge curves are smooth, bounded, and deterministic.
@@ 5.6.3 MIT-BIH Electrocardiogram Dataset
We use the MIT-BIH Arrhythmia dataset MITBIH_dataset_1980; MITBIH_url; physionet2000. ECG morphology (P, QRS, T waves) forms a low-variability compact manifold.
@@ 6.3.2 Synthetic ECG Generators
We consider the well-known low-dimensional ECG model proposed by McSharry et al. mcsharry2003dynamical, which generates ECG-like signals through a nonlinear dynamical system designed to approximate the P–QRS–T morphology. While the resulting waveforms are visually recognizable, they differ substantially from real ECG signals in fine temporal structure, smoothness, and relative amplitude balance.
@@ 9 Limitations and Future Work
By the Arzelà–Ascoli Theorem [rudin1991functional, royden2010real], any uniformly bounded and equicontinuous family of functions has compact closure in $C^{0}([0,T])$. Thus $\mathcal{M}$ is compact. ∎
@@
By the Universal Approximation Theorem [cybenko1989approximation, hornik1991approximation], for every $\varepsilon>0$ there exists a neural network $N_{\varepsilon}:\mathbb{R}^{N}\to\mathbb{R}$ such that
@@ Appendix D.2 McSharry dynamical ECG generator
The McSharry generator [mcsharry2003dynamical] models ECG signals using a low-dimensional nonlinear dynamical system designed to produce a stereotyped P–QRS–T morphology. While simplified, this model captures the gross temporal structure of cardiac cycles.

### v3

# The Blueprints of Intelligence A Functional–Topological Foundation for Perception and Representation

@@ 1 Introduction
Throughout this work, $\mathcal{M}$ denotes the set of realized signals generated by the physical system—the observed perceptual manifold—rather than the full sensory space. When the governing equations are unknown, both the manifold structure and its radius must be inferred directly from the stream of observations, leading naturally to a self-supervised formulation. This view is consistent with contemporary approaches to autonomous intelligence LeCun (2022), which emphasize that learning arises from discovering the set of admissible representations produced by the world rather than from external supervision.
@@
    We introduce a topological framework in which real-world perceptual sets are modeled as compact subsets of $C^{0}$ with finite Hausdorff radius and stable invariants Rudin (1991); Royden and Fitzpatrick (2010).
@@
    We show that deterministic physical processes induce continuous perceptual functionals that can be approximated by universal function approximators Cybenko (1989); Hornik (1991).
@@
2 Related Work
Joint Embedding Predictive Architectures (JEPA) show that useful world representations can be learned by predicting latent embeddings instead of raw data. Recent vision–language instantiations such as VL-JEPA predict continuous target text embeddings from visual inputs and textual queries, achieving competitive or superior performance to classical token-generative VLMs under comparable training conditions Chen et al. (2025). These approaches, however, do not provide a formal characterization of the geometry of the underlying signal manifolds that makes such prediction feasible, which is the focus of the present work.
@@ 3.2 Compactness of the perceptual set
This result follows directly from the Arzelà–Ascoli theorem Royden and Fitzpatrick (2010); Rudin (1991). Compactness implies that the system cannot produce arbitrary variability: all realized signals lie inside a bounded, closed, finite-variability geometric region in function space.
@@ 4.2 Universal approximation of perceptual mappings
Proof: See Appendix A. Follows from the Universal Approximation Theorem Cybenko (1989); Hornik (1991) and compactness of $\mathcal{M}$.
@@ 5.1 The perceptual radius
Using the Hausdorff metric $d_{H}$ Edgar (2008), the perceptual radius is defined as:
@@ 6.3 Distance metric and Hausdorff evaluation
In finite dimensions, all norms are equivalent (Rudin, 1991, Ch. 1), so compactness, convergence, and the existence of a finite perceptual radius are preserved under cosine distance, although the numerical value of the radius may differ between metrics.
@@ 6.6.1 Railway Point Machine Current Traces
We use the public Chinese Railway Point Machine dataset Li et al. (2020a, b). The signals exhibit a characteristic deterministic structure (inrush, plateau, closure peak), ideal for testing compactness and Hausdorff radius.
@@ 6.6.2 NASA Battery Aging Dataset
We use the NASA Ames Battery Dataset Saha and Goebel (2007a, b, 2011). Battery discharge curves are smooth, bounded, and deterministic.
@@ 6.6.3 MIT-BIH Electrocardiogram Dataset
We use the MIT-BIH Arrhythmia dataset Moody and Mark (2001, 1980); Goldberger et al. (2000). ECG morphology (P, QRS, T waves) forms a low-variability compact manifold.
@@ 7.3.2 Synthetic ECG Generators
We consider the well-known low-dimensional ECG model proposed by McSharry et al. McSharry et al. (2003), which generates ECG-like signals through a nonlinear dynamical system designed to approximate the P–QRS–T morphology. While the resulting waveforms are visually recognizable, they differ substantially from real ECG signals in fine temporal structure, smoothness, and relative amplitude balance.
@@ 8.1 Implications for perception and machine learning
These ideas are not limited to time-series signals. Public trademark offices such as the EUIPOEuropean Union Intellectual Property Office (EUIPO) (2019) deploy AI-based logo similarity tools that embed marks into a perceptual representation space and retrieve visually similar candidates from large registers. Conceptually, admissible logos for a given semantic message occupy a compact region of this space, while marks that are “too different” fall outside the corresponding perceptual manifold. While implementation details differ, our framework provides a natural geometric interpretation of such systems: similarity search can be seen as testing whether a query logo lies within a finite-radius neighborhood of an existing manifold of admissible realizations.
@@ 8.4 Industrial impact and practical implications
8.5 Relation to JEPA and latent prediction models
Our functional–topological framework is complementary to JEPA-based architectures such as VL-JEPA Chen et al. (2025). While VL-JEPA empirically shows that predicting target embeddings in a continuous latent space leads to efficient vision–language learning, our results explain why such an approach is plausible in real-world domains: deterministic physical processes generate compact perceptual manifolds with finite Hausdorff radius, on which continuous perceptual functionals are universally approximable. An interesting direction for future work is to use manifold-level quantities such as empirical Hausdorff radius and saturation behavior to analyze or regularize the embedding spaces learned by JEPA-style models like VL-JEPA Chen et al. (2025).
@@ 10 Limitations and Future Work
By the Arzelà–Ascoli Theorem [14, 13], any uniformly bounded and equicontinuous family of functions has compact closure in $C^{0}([0,T])$. Thus $\mathcal{M}$ is compact. ∎
@@
By the Universal Approximation Theorem [2, 6], for every $\varepsilon>0$ there exists a neural network $N_{\varepsilon}:\mathbb{R}^{N}\to\mathbb{R}$ such that
@@ Appendix D.2 McSharry dynamical ECG generator
The McSharry generator [10] models ECG signals using a low-dimensional nonlinear dynamical system designed to produce a stereotyped P–QRS–T morphology. While simplified, this model captures the gross temporal structure of cardiac cycles.